OpenStudy (anonymous):

Does the line (x-7)/3=(y+8)/6=(z+4)/5 and the point P(10,-2,1) determine a plane? How would I answer this?

OpenStudy (anonymous):

\[\frac{ x-7 }{ 3 }= \frac{ y+8 }{ 6 }=\frac{ z+4 }{ 5 }\]

OpenStudy (rational):

does the given point lie on the line ?

OpenStudy (anonymous):

it doesn't say, but i think it does

OpenStudy (rational):

simply plugin the given point into the given equaiton of line and see if it satisfies

OpenStudy (anonymous):

ok I did and they all equal to each other. That means it does determine a plane?

OpenStudy (rational):

Nope that means the given point lies on the given line

OpenStudy (rational):

so the line and point just form a line they dont realize a plane

OpenStudy (rational):

you need 3 noncollinear points for a plane

OpenStudy (anonymous):

oh ok thanks!

OpenStudy (rational):

yw